OpenStudy (anonymous):

Find a power series representation for the function, and find the interval of convergence: 1/(1-x^3)

OpenStudy (anonymous):

looks just like \[\frac{1}{1-x}=\sum x^n\] only with \(x^3\) instead of \(x\)

OpenStudy (anonymous):

replace \(x\) by \(x^3\) and you are done

OpenStudy (anonymous):

btw same radius of convergence, since \(|x|<1\iff |x^3|<1\)

OpenStudy (anonymous):

Wait so the answer is the same thing? There are no major steps? Can you lead me through it...?

OpenStudy (anonymous):

@satellite73

OpenStudy (kirbykirby):

There really isn't much other than to replace \(x\) by \(x^3\) as satellite said. In their first reply, they gave you the series representation for 1/(1-x). If you just change the \(x\) to \(x^3\), then you just replace \(x\) with \(x^3\) on the right hand side: \(\large \sum (x^3)^n\). The idea is to see if your given series can "look" like the basic series you know for \(1/(1-x)\).

OpenStudy (kirbykirby):

It may seem a bit trivial, but maybe it will make more sense as you do more of these types of problems.